### v6

# Over-Squashing in Graph Neural Networks: A Comprehensive survey

Abstract
Graph Neural Networks (GNNs) revolutionize machine learning for graph-structured data, effectively capturing complex relationships. They disseminate information through interconnected nodes, but long-range interactions face challenges known as ”over-squashing”. This survey delves into the challenge of over-squashing in GNNs, where long-range information dissemination is hindered, impacting tasks reliant on intricate long-distance interactions. It comprehensively explores the causes, consequences, and mitigation strategies for over-squashing. Various methodologies are reviewed, including graph rewiring, novel normalization, spectral analysis, and curvature-based strategies, with a focus on their trade-offs and effectiveness. The survey also discusses the interplay between over-squashing and other GNN limitations, such as over-smoothing, and provides a taxonomy of models designed to address these issues in node and graph-level tasks. Benchmark datasets for performance evaluation are also detailed, making this survey a valuable resource for researchers and practitioners in the GNN field.
Index Terms:
Graph Neural Networks (GNNs), Over-squashing, Over-smoothing, Graph-rewiring, Graph transformers
I Introduction
In recent years, the explosion of data in various domains has led to an increased interest in harnessing the power of graph structures for modeling complex relationships [1, 2, 3, 4, 5]. Graphs, which consist of nodes and edges representing entities and their connections, respectively, have emerged as a fundamental data representation in fields such as social networks [2, 6, 7], recommendation systems[8, 9, 10, 11], biology [12, 13], and more. As the diversity and complexity of graph-structured data grow, so does the demand for advanced tools to analyze and understand these intricate relationships.
The burgeoning interest in leveraging graph-structured data has given rise to a remarkable class of machine learning models known as Graph Neural Networks (GNNs) [14, 15, 16]. GNNs represent a novel approach to learning representations, allowing the unified capture of both local and global information of nodes within graphs [17, 18]. In essence, GNNs extend traditional neural network architectures to accommodate graph data, where nodes symbolize entities and edges signify relationships. This extension has paved the way for a myriad of applications, spanning from node classification [19, 20, 21] and link prediction to broader graph-level tasks such as community detection [6, 22] and molecular property prediction [23, 24]. By leveraging the underlying graph structure, GNNs facilitate information propagation and aggregation, allowing them to capture intricate patterns that traditional machine learning models find challenging to discern.
Notwithstanding their remarkable achievements, GNNs are not immune to certain inherent limitations, including over-smoothing [25, 26], vanishing gradients [27, 28], Out-of-Distribution (OOD) data challenges [29, 30], overfitting [31], and the relatively less explored phenomenon of over-squashing [32, 33, 34]. While exhaustive research has been dedicated to addressing the former issues, the latter—over-squashing—remains relatively less explored.
Over-squashing is a phenomenon that manifests in tasks requiring the integration of information from distant nodes[32, 35], primarily through edges that serve as bottlenecks within graph data. To put it succinctly, over-squashing denotes the distortion-prone nature of information transfer between nodes that are widely separated[34]. This distortion emerges due to the inherent tension between the limited feature representation capacity of graph embeddings and the exponential growth in the number of neighbors as graphs expand. This interplay often hampers the faithful transmission of distant information.
This survey article aims to provide a comprehensive panorama of this specific limitation. We delve into the intricate nuances of over-squashing, shedding light on its conceptual framework and its implications. Additionally, we meticulously outline the repertoire of methods proposed thus far to grapple with this intricate issue. By presenting a systematic exploration of the landscape, we contribute to a deeper understanding of over-squashing’s impact on GNNs and offer insights into the evolving strategies engineered to surmount this challenge.
To summarize, this paper makes the following key contributions:
  1. 1.
    Pioneering Survey: This paper serves as the inaugural comprehensive survey on ’over-squashing,’ a pivotal limitation in message-passing graph neural networks. It addresses a burgeoning area of interest among researchers.
  2. 2.
    Systematic Categorization: We provide a systematic categorization of existing methods, offering a detailed taxonomy that simplifies the understanding of various strategies to mitigate over-squashing.
  3. 3.
    Benchmark Datasets: We extensively discuss commonly used benchmark datasets employed for evaluating models in the context of over-squashing, both at the node and graph levels.
  4. 4.
    Added Value: Additionally, this survey explores the interplay of over-squashing with other fundamental GNN limitations, such as ’over-smoothing,’ providing a more holistic perspective on the challenges faced in this domain.
This article is organized as follows: In Section II, we provide a comprehensive background on graph neural networks (GNNs) (II-A), detailing the concept of over-squashing and its mathematical underpinnings. Section III delves into popular methods for mitigating over-squashing in GNNs, including spatial graph rewiring (III-A) and spectral graph rewiring (III-B). In Section IV, we explore the intricate interplay between over-squashing and over-smoothing in GNNs, analyzing their trade-offs and presenting unified approaches to address both challenges simultaneously. Section V examines novel techniques beyond rewiring methods that tackle over-squashing. Furthermore, Section VI introduces commonly used benchmark datasets for evaluating over-squashing in GNNs, applicable to both node classification and graph classification tasks. Finally, Section VII offers concluding remarks and outlines promising avenues for future research, providing insights into underexplored directions for further exploration.
II Background
II-A Graph Neural Networks
A Graph Neural Network is a specialized neural network architecture designed for processing data structured as graphs. The central concept behind GNNs involves the iterative collection of information from neighboring nodes and the updating of node features through multiple layers.
Let’s delve deeper into this concept. Imagine a graph, denoted as $G=(V,E)$, where $V$ represents the set of nodes with $|V|=n$ and $E$ signifies the set of edges (or links), adhering to $E\subseteq V\times V$. If for $u,v\in V$ is connected by an edge, we write it as $(u,v)\in E$. The adjacency matrix $A$ is defined by $A_{uv}=1$ if $(u,v)\in E$, and $A_{ij}=0$ otherwise. Within this framework, each node $v\in V$ comes with an associated feature vector $x_{v}\in\mathbb{R}^{d_{0}}$ effectively encapsulating information pertaining to the attributes of the node $v$. We use $X\in\mathbb{R}^{n\times d_{0}}$ as feature matrix. In the realm of GNNs, the primary goal is to acquire effective representations for nodes, links, and even entire graphs. This is achieved through a fundamental process known as message-passing, initially defined by Gilmer et al. [4] and further elaborated by Zhang et al. [36].
In this message-passing process, GNNs iteratively enhance node representations using the following equations:
At layer $l$:
$h^{(l)}_{u}$ $=UP_{l}\biggl{\{}h^{(l-1)}_{u},\vspace{1cm}AGG_{l}\{h^{(l-1)}_{v}\mbox{ where }v\in N_{u}\}\biggr{\}}$ (1)
Here, $h^{(l-1)}_{u}$ represents the node representation at the $(l-1)$-st layer, typically initialized with the node’s feature at the initial layer $h^{(0)}_{u}=x_{u}$. $N_{u}$ denotes the set of $1$-hop neighboring nodes of node $u$:
$N_{u}=\{v\in V:(v,u)\in E\}.$
The aggregation function family, denoted as $AGG_{l}(\cdot)$, is responsible for mixing the information from neighboring nodes and has the form:
$AGG_{l}:\mathbb{R}^{d_{(l-1)}}\times\mathbb{R}^{d_{(l-1)}}\to\mathbb{R}^{d_{(l-1)}^{\prime}}$
Furthermore, the update function family, referred to as $UP_{l}(\cdot)$, integrates the aggregated information into the node’s representation and has the form:
$UP_{l}:\mathbb{R}^{d_{(l-1)}}\times\mathbb{R}^{d_{(l-1)}^{\prime}}\to\mathbb{R}^{d_{(l)}}$
In the context of message passing GNNs, $AGG$ and $UP$ typically serve as attention mechanisms and activation functions, respectively. Through iterative application of this message-passing mechanism, GNNs progressively enhance node representations by accounting for their associations with neighboring nodes. This iterative refinement process plays a vital role in capturing both structural and semantic insights embedded within the graph.
II-B Over-squashing
The issue of over-squashing is a recognized challenge encountered in Message Passing Neural Networks (MPNNs) when messages traverse distant nodes. This problem arises due to the rapid expansion of a node’s receptive field, leading to the compression of numerous messages into fixed-size vectors. Topping et al. [35] have formally substantiated this phenomenon through a sensitivity analysis of the Jacobian matrix of node features.
To elaborate further, let’s consider a receptive field denoted as $B_{r}=\{v\in V:d_{G}(u,v)\leq r\}$ associated with an $r$-layer GNN. Here, $d_{G}$ represents the shortest-path distance between node $u$ and node $v$, and $r$ is a natural number. The shortest-path between node $u$ and $v$ is denoted by $P_{uv}$. The Jacobian matrix $\frac{\partial h^{(r)}_{u}}{\partial x_{v}}$ quantifies the sensitivity of a node representation $h^{(r)}_{u}$ to a specific input feature $x_{v}$ in node $v$. Over-squashing can be conceptualized as the incapacity of $h^{(r)}_{u}$ to be influenced by $x_{v}$ at a distance of $r$. Topping et al. [35] have established the following bounds to quantify the effect of over-squashing:
$\frac{\partial h^{(r+1)}_{u}}{\partial x_{v}}\leq(\alpha\beta)^{r+1}(\tilde{A}^{r+1})_{uv}.$ (2)
In equation (2), $\tilde{A}=D^{-\frac{1}{2}}(A+I)D^{-\frac{1}{2}}$ represents the normalized adjacency matrix, where $D_{ii}=\sum_{j=1}^{n}A_{ij}$ stands for the degree matrix of $(A+I)$. Note that ${A}_{ij}$ denotes the entry in row $i$ and column $j$ of matrix ${A}$. Inequality (2) conveys that if $|\nabla UP_{l}|\leq\alpha$ and $|\nabla AGG_{l}|\leq\beta$ for $10\leq l\leq r+$ 1, where $\nabla g$ denotes the Jacobian of a map $g$, the propagation of messages can be controlled by an appropriate power of $\tilde{A}$. Specifically, this inequality emphasizes how the influence of input features diminishes exponentially as the distance $r$ increases, which becomes particularly pronounced when the size of $B_{r}$ grows exponentially.
On the same assumption as for bound (2), Black et al. [37] provide the following bound:
$\frac{\partial h^{(r+1)}_{u}}{\partial x_{v}}\leq(2\alpha\beta)\sum_{l=0}^{r+1}(\tilde{A}^{l})_{uv}.$ (3)
The primary distinction between bounds (2) and (3) lies in their conditions: the former holds when the vertices $u$ and $v$ must be precisely at a distance $r+1$ from each other, while the latter applies to any pair of vertices.
Black et al. [37] additionally established a bound to quantify the over-squashing effect, which correlates heightened effective resistance between node pairs to the sensitivity of a node representation $h^{(r)}_{u}$ to a specific input feature $x_{v}$.
Definition II.1
The effective resistance between two nodes $u$ and $v$ is defined as:
$R_{u,v}=\left(\frac{1}{\sqrt{d_{u}}}\mathbf{1}_{u}-\frac{1}{\sqrt{d_{v}}}\mathbf{1}_{v}\right)^{T}\tilde{L}^{+}\left(\frac{1}{\sqrt{d_{u}}}\mathbf{1}_{u}-\frac{1}{\sqrt{d_{v}}}\mathbf{1}_{v}\right),$ (4)
where $\mathbf{1}_{v}$ is the indicator vector of vertex $v$, $d_{v}$ denotes the degree of vertex $v$, and $\tilde{L}^{+}$ is the pseudoinverse of the normalized Laplacian matrix $\tilde{L}=I-\tilde{A}$.
The effective resistance quantifies the level of connectivity between vertices $u$ and $v$ within the graph $G$. Essentially, when multiple paths exist between two nodes, the effective resistance $R_{u,v}$ tends to be small, signifying higher connectivity. Conversely, when the available paths between two nodes are limited, $R_{u,v}$ becomes larger, indicating lower connectivity.
Based on the effective resistance between nodes $u$ and $v$, Black et al [37] provide the following bound to quantify the impact of each node representation $h^{(r)}_{u}$ with respect to $x_{v}$.
Theorem II.1
For a connected graph $G$, let $u,v\in V$ and $\|\nabla UP_{l}\|\leq\alpha$ and $\max\{\|\nabla AGG_{l}\|,1\}\leq\beta$ for all $l=0,\ldots,r$. Let $d_{\text{max}}$ and $d_{\text{min}}$ be the maximum and minimum degrees of nodes $u$ and $v$, and $\max\{|\mu_{2}|,|\mu_{n}|\}\leq\mu$ where $\mu_{n}\leq\mu_{n-1}\leq\cdots\leq\mu_{1}=1$ denote the eigenvalues of $\tilde{A}$. Then,
$\left\|\frac{\partial h^{(r)}_{u}}{\partial x_{v}}\right\|\leq(2\alpha\beta)^{r}\frac{d_{\text{max}}}{2}\left(\frac{2}{d_{\text{min}}}\left(r+1+\frac{\mu^{r+1}}{1-\mu}\right)-R_{u,v}\right).$ (5)
Theorem II.1 implies that vertices with lower effective resistance exert a stronger influence on each other during message passing. Specifically, the node feature $h^{(r)}_{u}$ at node $u$ in layer $r$ is more heavily influenced by the initial node feature $x_{v}$ at node $v$. Hence, it characterizes the over-squashing effect in GNN as proportional to the effective resistance between nodes of the graph. This notion aligns with intuition, as effective resistance reflects the connectivity and number of paths between $u$ and $v$. When there are numerous short paths connecting $u$ and $v$, the effective resistance between them decreases, facilitating stronger interaction hence low over-squashing. In a parallel vein, Di Giovanni et al. [34] have undertaken similar methodologies, ultimately converging on a shared conclusion. Their findings underline the pivotal role of effective resistance in influencing the degree of over-squashing within GNNs. They delve into the impact of GNN architecture’s width and depth on the occurrence of over-squashing. In their work, Di Giovanni et al. [34] build upon their previous findings and concentrate on two pivotal factors: the network’s architecture, characterized by weight norms and depth, and the intrinsic graph structure, evaluated using commute times. Commuting time, in the context of a graph, refers to the expected number of steps it takes for a random walk to travel between two specific nodes. It’s like measuring how long it would typically take for someone to move from one node to another and coming back by randomly traversing the edges of the graph. It is well known that [38] commute time between node $u$ and $v$ (denoted as $Com(u,v)$) is directly proportional to effective resistance $R_{u,v}$. In mathematical terms, this relationship is expressed as $Com(u,v)=2|E|R_{u,v}$, with $E$ representing the set of edges in the graph.
In [34], the quantification of the influence of $x_{v}$ on the node representation of node $u$ at any layer $l<r$ is refined as the symmetric Jacobian obstruction between nodes $u$ and $v$. This refinement is achieved by defining
$O^{(r)}_{u,v}=\sum_{l=0}^{r}\left\|J^{(r)}_{l}(u,v)\right\|,$ (6)
[…]
$J^{(r)}_{l}(u,v)=\frac{1}{d_{u}}\frac{\partial h^{(r)}_{u}}{\partial h^{(l)}_{u}}-\frac{1}{\sqrt{d_{u}d_{v}}}\frac{\partial h^{(r)}_{u}}{\partial h^{(l)}_{v}}+\frac{1}{d_{v}}\frac{\partial h^{(r)}_{v}}{\partial h^{(l)}_{v}}-\frac{1}{\sqrt{d_{u}d_{v}}}\frac{\partial h^{(r)}_{v}}{\partial h^{(l)}_{u}}$
Intuitively, $J^{(r)}_{l}(u,v)$ reflects the sensitivity between nodes $u$ and $v$. When nodes $u$ and $v$ are less sensitive to each other, $J^{(r)}_{l}(u,v)$ tends to be larger; conversely, it is smaller when the communication is robust. $O^{(r)}_{u,v}$ quantifies the influence of the initial feature of node $v$ on $h^{(r)}_{u}$ and can be bounded by the commute time.
Theorem II.2
For a specific MPNN as in (1) with $AGG_{l}=W^{(l)}$. Let $\nu$ be the minimal singular value and $w$ be the maximal spectral norm of the matrix $W^{(l)}$. Assuming all paths in the MPNN graph are activated with success probability $\rho$ then there exists a constant $\epsilon_{G}$ independent of $u$ and $v$, such that:
$\epsilon_{G}(1-o(l))\frac{\rho}{\nu}\frac{Com(u,v)}{2|E|}\leq O^{(l)}_{u,v}\leq\frac{\rho}{w}\frac{Com(u,v)}{2|E|}$
where $o(l)$ approaches 00 exponentially fast as $l$ increases.
Previously discussed, a smaller value of $O^{(r)}_{u,v}$ indicates that node $u$ is more sensitive to $v$ in the MPNN, and vice versa. Therefore, Theorem II.2 suggests that nodes with shorter commute times will exchange information more effectively in an MPNN, while conversely, nodes with longer commute times will exchange information less effectively. Consequently, they infer that over-squashing becomes problematic when the task relies on interactions between nodes with high commute times.
In previous works [35, 37, 34], researchers have formalized the phenomenon of over-squashing by examining how the Jacobian of node features is influenced by topological properties of the graph, such as curvature, effective resistance, or commute time. However, in [39], the authors conducted a more comprehensive analysis and introduced a novel over-squashing metric. This metric takes into account various aspects including GNN architecture, parameters, graph topology, and downstream tasks. As part of their investigation, they established upper bounds on the capacity of MPNNs, defined by the pair $(r,w)$, where $r$ represents the number of layers and $w$ denotes the maximum spectral norm of the weights. A higher capacity, achieved by increasing either $r$ or $w$, or both, indicates greater expressive capability of the MPNN, allowing for more extensive mixing among node features. Importantly, they introduced the concept of ”over-squashing,” which is intricately linked to the maximum node mixing capacity of MPNNs and operates inversely to it.
Definition II.2
Consider a (twice differentiable) function $y_{G}$ (in the context of MPNN it serves as ground truth function for graph-level tasks). The maximal mixing induced by $y_{G}$ among variables $x_{u}$ and $x_{v}$, where $u\neq v$, is given by:
$\text{mix}_{y_{G}}(u,v)=\max_{x_{i}}\max_{1\leq\alpha,\beta\leq d_{0}}\left|\frac{\partial^{2}y_{G}(X)}{\partial x_{\alpha}(u)\partial x_{\beta}(v)}\right|,$
where $x_{u}(i)$ denotes the $i$th component in the vector $x_{u}$.
Definition II.3
In the context of an MPNN comprising $r$ layers and $w$ representing the maximum spectral norm of its weights, we define the tuple $(r,w)$ as indicative of the MPNN’s capacity.
The over-squashing of nodes $v$ and $u$ in an MPNN, characterized by its capacity $(m,w)$, is quantified by the function $OSQ_{v,u}(m,w)$. For a given MPNN with capacity $(r,w)$, the over-squashing effect between nodes $u$ and $v$ can be characterized by the pairwise mixing induced by an MPNN with the same capacity $(r,w)$. Specifically, higher over-squashing between $u$ and $v$ corresponds to lower mixing. This relationship can be expressed as:
$OSQ_{u,v}(r,w)<\left(\text{mix}_{y_{G}}(u,v)\right)^{-1}$
Hence, an MPNN operating on a graph $G$ with node features $X$ may exhibit over-squashing depending on the complexity of the task it’s designed to solve. This phenomenon varies based on the nodes $v$ and $u$ involved, being more pronounced for nodes with greater separation or higher commute time, highlighting the significant influence of graph topology on over-squashing.
III Handling over-squashing in GNNs
In scenarios where tasks necessitate spanning multiple layers within a network, the depth of the network often mirrors the range of interactions between nodes. Nevertheless, a rising number of layers corresponds to an exponential increase in the number of nodes contributing to the receptive field of each individual node. This amplification leads to the phenomenon of over-squashing [32, 35]. Essentially, over-squashing manifests as a compression of information originating from a receptive field that encompasses numerous nodes. This compression results in fixed-length node vectors, impeding the accurate propagation of messages from distant nodes. This distortion takes shape due to graph bottlenecks that emerge as the number of $k$-hop neighbors undergoes exponential growth with each $k$.
In a bid to surmount these challenges, the literature has proposed strategies such as graph rewiring [32, 35, 40] and pooling[41, 42, 43]. Graph rewiring is a technique used to modify the structure of a graph, particularly the edges, with the goal of enhancing the performance of GNNs. Essentially, rewiring entails applying an operation $R$ to $G=(V,E)$, resulting in a new graph $R(G)=(V,R(E))$ with modified connectivity while retaining the same set of vertices. We extend the MPNN framework (1) to incorporate the rewiring operation $R$ as follows:
$h^{(l)}_{u}=\text{UP}_{l}\left(h^{(l-1)}_{u},{AGG_{(l)}}_{G}\left(\{h^{(l-1)}_{v}:(u,v)\in E\}\right),\right.\\ \left.{AGG_{(l)}}_{R(G)}\left(\{h^{(l-1)}_{v}:(u,v)\in R(E)\}\right)\right),$ (7)
where node features are updated based on the information gathered from both the original graph $G$ and the rewired graph $R(G)$ using potentially independent aggregation maps. Many GNN models based on rewiring simply exchange messages over $R(G)$, i.e., they utilize ${AGG_{(l)}}_{G}=0$.
Graph rewiring typically involves making adjustments to the edges, either by adding new edges [44, 37], removing existing ones [45], a combination of both[33] or replacing existing ones by new ones [35, 46]. The primary objective of graph rewiring is to optimize the flow of information within the graph, making it more suitable for specific tasks like graph or node classification, as well as link prediction. Many GNNs implicitly incorporate the concept of graph rewiring, which involves various techniques such as utilizing Cayley graphs [47] or introducing virtual nodes [48]. Some research explores directly modifying the graph connectivity to address noise [49] or facilitate multi-hop aggregations [50].
The challenge of over-squashing within GNNs has spurred the development of various methodologies, each aiming to alleviate this phenomenon. Broadly, these methods can be categorized into two types of graph rewiring methods, each offering unique insights into the resolution of the over-squashing predicament.
III-A Spatial Graph Rewiring Methods:
Spatial graph rewiring involves modifying the edges of a graph by considering the geometric or spatial relationships between nodes. By optimizing the spatial connections, the rewiring process seeks to improve the overall performance of GNNs in tasks that involve understanding spatial dependencies and relationships among graph elements. The following spatial graph rewiring techniques have been proposed in the literature to mitigate the over-squashing issue, particularly in tasks involving long-range dependencies.
Alon and Yahav [32] were among the first researchers to address the challenge of over-squashing within GNNs. They pointed out that, particularly in prediction scenarios relying on extensive node interactions, GNNs often struggle with over-squashing. This challenge arises because the number of nodes in each node’s receptive field grows exponentially with an increasing number of layers in the network. To address this issue, they proposed a straightforward solution: introducing a fully-adjacent matrix in the final GNN layer. In this method, called the Fully-Adjacent (FA) layer, the GNN layer is reshaped to connect every pair of nodes with an edge. Specifically, in a GNN with $r$ layers, modifications are made only to the $r$-th layer by replacing it with an FA layer. The FA layer ensures that every pair of nodes is directly connected by an edge. Converting an existing layer (1) to be fully-adjacent means that for every node $u\in V$, $N_{u}:=V$ in that layer only. At layer $r$, the transformation can be expressed as:
$h^{(r)}_{u}$ $=UP_{r}\left\{h^{(r-1)}_{u},AGG_{r}\{h^{(r-1)}_{v}\text{ where }v\in V\}\right\}.$ (8)
This transformation does not change the type of layer or add weights but only alters the adjacency of a data sample in a single layer. As a result, the $r-1$ graph layers can still exploit the graph structure using their original sparse topology, while only the $r$-th layer becomes an FA layer, enabling direct interactions among node representations that are aware of the network’s topology and facilitating consideration of nodes beyond their initial neighbors. Topping et al. [35] introduced a novel approach that utilizes graph curvature to address the over-squashing problem in GNNs. Their method, known as Stochastic Discrete Ricci Flow (SDRF), aims to surgically modify negatively-curved edges in order to reduce bottlenecks without significantly altering the statistical properties of the input graph. The core idea of SDRF is to target edges with high negative curvature, which are identified as key contributors to over-squashing. The SDRF algorithm operates in two main steps. First, it identifies edges with minimal Ricci-curvature [^1] Given a graph $G=(V,E)$ with vertices $V$ and edges $E$, the Ricci curvature of an edge $e=(u,v)$ is denoted as $\text{Ric}_{e}$ and is defined in terms of the ratio of the effective resistance between the endpoints of the edge and the length of the edge itself. It can be calculated using various methods, such as discrete Laplacian operators or spectral methods, for details see [51]., indicating potential over-squashing issues. Then, it constructs additional supportive edges around these identified edges to reinforce their structural context within the graph. By strengthening these edges, the algorithm enhances the information flow within the GNN and mitigates the adverse effects of over-squashing.
Black et al. [37] conducted an in-depth analysis of over-squashing by exploring its connection with effective resistance between node pairs, as discussed previously in Section II-B. They proposed a novel graph rewiring method known as the Greedy Total Resistance (GTR) technique to address this issue. The GTR technique aims to minimize the effective resistance between nodes by strategically rewiring the graph. They derived a formula quantifying the impact of adding a specific edge $(u,v)$ on the reduction of total resistance. The total change in resistance when adding an edge $(u,v)$ to a connected graph $G$ with $n$ vertices can be expressed as the difference between the total resistance before and after the addition:
$R_{\text{tot}}(G)-R_{\text{tot}}(G\cup\{u,v\})=n\cdot\frac{B_{u,v}^{2}}{1+R_{u,v}}$
Here, $n$ represents the number of vertices in the graph, $B_{u,v}$ is the biharmonic distance of the node pair defined as:
$B_{u,v}=\sqrt{\left(\mathbf{1}_{u}-\mathbf{1}_{v}\right)^{T}\left(\tilde{L}^{+}\right)^{2}\left(\mathbf{1}_{u}-\mathbf{1}_{v}\right)}$
and $R_{u,v}$ denotes the effective resistance, as given in Equation (4). The GTR method operates on a greedy principle, strategically adding edges to the graph to maximize $\frac{B_{u,v}^{2}}{1+R_{u,v}}$ and consequently minimize total resistance.
Topping et al. [35] delved into over-squashing, identifying it as the presence of edges with negative curvature. On the other hand, Black et al. and Di Giovanni [37, 34] linked over-squashing to effective resistance or commute time between nodes. In contrast, Gabrielsson et al. [52] proposed a distinct rewiring technique leveraging transformer-inspired positional encoding within a graph framework to expand the receptive field of each node in GNNs. Their method, unlike others, is model-agnostic and focuses on augmenting the original graph with additional nodes and edges using positional encoding as node and/or edge features. This augmentation process does not involve removing existing edges or nodes, ensuring that the original graph remains a subset of the new rewired graph. Specifically, given a graph $G$ and a positive integer $r$, they add edges to all nodes within $r$-hops in $G$ to create a new graph $G_{r}$. Additionally, a fully-connected CLS (classification) node is introduced in $G_{r}$, which connects to all nodes in the graph. However, there’s a concern that in the case of large $r$, the connectivity in the rewired graph $G_{r}$ may become too dense, potentially leading to the loss of the original graph’s topology, a crucial cue for graph-based learning. To address this, they encode the original topology of $G$ into the rewired graph $G_{r}$ using positional encodings [^2] A Positional Encoding refers to a feature that characterizes the spatial arrangement of nodes within a graph, providing information about their global or local positions. These features are often associated with random walk measures and the eigenvectors of the Laplacian matrix [53]., which serve as node and/or edge features. Another approach that probabilistically rewires the graph based on the given prediction task is the Probabilistically Rewired Message-Passing Graph Neural Network (PR-MPNN) [54]. PR-MPNN leverages recent advancements in exact and differentiable k-subset sampling [55] to probabilistically rewire the graph during the learning process. By doing so, PR-MPNN learns to add edges that are relevant to the prediction task while omitting less beneficial ones, effectively addressing issues like over-squashing and under-reaching. Gutteridge et al. [56] propose a novel layer-dependent rewiring technique to address challenges posed by existing rewiring approaches [57, 52] in GNNs. They highlight that some rewiring methods, while aiming to improve connectivity for long-range tasks, compromise the inductive bias provided by graph distance. These approaches enable instant communication between distant nodes at every layer, disrupting the inherent structure. To overcome this, Gutteridge and colleagues introduce a layer-dependent rewiring strategy, dynamically rewired message-passing with delay (DRew) that gradually densifies the graph. This technique allows for enhanced connectivity while preserving the inductive bias provided by graph distance. Additionally, they incorporate a delay mechanism that facilitates skip connections based on both node distance and layer. This mechanism ensures that the graph’s inductive bias is retained, providing a more nuanced and context-aware rewiring approach for improved performance in long-range tasks
TABLE I: Taxonomy of Methods for Mitigating Limitations in GNNs Stemming from Long-Range Propagation. We categorize these approaches into two primary domains: spectral and spatial methods. Spectral methods in GNNs harness graph eigenvalues and eigenvectors for signal analysis and propagation, while spatial methods involve direct processing of node and edge information to capture local graph structures.
Category Approach Methods Task Targets Graph Type
Nodes Graphs Over-smoothing Over-squashing Homophily Heterophily
Spatial Method Fully adjacent Rewiring FA [32]
Spatial Method Curvature-based rewiring SDRF [35]
Spatial Method Total Resistance GTR [37]
Spatial Method Probabilistic rewiring PR-MPNN [54]
Spatial Method Layer-dependent Rewiring DRew [56]
Spatial Method Training-free reservoir model GESN [58]
Spatial Method Curvature-based rewiring BORF [40]
Spectral Method Expander graph EGP [47]
Spectral Method Commute time based rewiring DiffWire [59]
Spectral Method Curvature-based Edge flip G-RLEF [46]
Spectral Method Edge addition FoSR [44]
Spectral Method Edge addition & removal SJLR [33]
Spectral Method Curvature-based edge drop CurvDrop [45]
Other Method Directional Propagation DGN[60]
Other Method Heat kernel as filter MHKG [61]
Other Method Graph ODEs A-DGN [62]
Other Method Curvature-based pooling CurvPool [43]
Other Method Topological information integration PASTEL [63]
Other Method Structural information propagation RFGNN [64]
III-B Spectral Graph Rewiring Methods:
To explain graph rewiring in the context of spectrum of the graph we would like to explain the connectedness of a graph with eigen values of the graph Laplacian. The connectedness of a graph $G$ can be measured via a quantity known as the Cheeger constant, denoted as $h(G)$, is defined as follows [65]:
$h(G)=\min_{(U\subset V)}\frac{\big{|}\{(u,v)\in E:u\in U,v\in V\setminus U\}\big{|}}{\min\big{(}vol(U),vol(V\setminus U)\big{)}}$
Here, $vol(U)$ represents the volume of set $U$ and is calculated as the sum of degrees of nodes $u\in U$.
The Cheeger constant, $h(G)$, essentially quantifies the energy required to divide graph $G$ into two separate communities. A smaller $h(G)$ implies that $G$ tends to have two communities with only a few connecting edges. In such cases, over-squashing is more likely to occur when information needs to traverse from one community to another. It’s important to note that while computing $h(G)$ is generally a complex task, the Cheeger inequality provides a useful relationship: $h(G)$ is approximately proportional to the smallest positive eigenvalue $\lambda_{1}$ of the graph Laplacian (spectral gap of the graph [65]). We are assuming $0=\lambda_{0}\leq\lambda_{1}\leq\cdots\lambda_{n}$ are the eigenvalues of the normalized graph Laplacian matrix ${L}$.
In light of this relationship, some recent approaches have proposed selecting a rewiring strategy that depends on the spectrum of $G$. The goal is to generate a new graph $R(G)$ that satisfies $h(R(G))>h(G)$. This strategy has been explored in the works [59, 47, 44]. The underlying assumption is that propagating messages over the rewired graph $R(G)$ can mitigate over-squashing. However, it’s important to note that this claim lacks formal analytical proof at this stage.
Deac et al. [47] proposed a model for graph classification task where propagation of information takes place on expander graphs in MPNN and named the method as Expander Graph Propagation (EGP) model. Deac argue that there are many tasks (being graph classification one of them) where only local node level interaction is not sufficient to predict the correct label, specifically for graph classification task where together with local interaction among node feature, global context of graph structure is also important. For such tasks one need to update node features in a manner that is mindful of the global properties of the graph just stacking more message passing layers on the input graph will not sufficient. EGP model strategically employs expander graphs to address challenges associated with bottleneck-induced issues in global information propagation within a graph. Expander graphs are family of graph structures that are fundamentally sparse $(|E|=O|V|)$, while having low diameter: thus, any two nodes in an expander graph may reach each other in a short number of hops, eliminating bottlenecks and oversquashing.
Definition III.1
A family $\{G_{i}\}$ of finite connected graphs is termed an expander family if there exists a constant $c\geq 0$ such that for every $G_{i}$ in the family, the first non-zero eigenvalue $\lambda_{1}(G_{i})$ satisfies $\lambda_{1}(G_{i})\geq c$.
Expander graphs can be interpreted in terms of Cheegers constant as: [66]
Theorem III.1
Consider an infinite collection $\{G_{i}\}$ of expander graphs with a uniform upper bound on their vertex degrees. Then, there exists a constant $\epsilon>0$ such that for all graphs in the collection, $h(G_{i})\geq\epsilon$.
Since expander graphs have higher Cheeger constants and will hence experience less severe problems arising due to bottleneck edges They proposed a method to efficiently construct expander of vertex approximately equal to $|V|$ using special linear group.
Definition III.2
For any positive integer $n$, the special linear group $\text{SL}(2,\mathbb{Z}_{n})$ represents the group of $2\times 2$ matrices whose entries are integers modulo $n$ and with determinant 1. One of its generating sets is:
$S_{n}=\left\{\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\begin{pmatrix}1&0\\ 1&1\end{pmatrix}\right\}.$
Theorem III.2
The Cayley graph family $\text{Cay}(SL(2,\mathbb{Z}_{n});S_{n})$ constitutes an expander family [67] with vertices equal to the number of elements in $SL(2,\mathbb{Z}_{n})$.
Once the expander graph with appropriate $n$ is constructed, EGP model follow a simple approach by running standard GNN layer on input structure i.e. $X,A$ followed by a GNN layer over the relevant Cayley graph. Suppose we define $A_{\text{Cay}}(n)$ as an adjacency matrix obtained from the Cayley graph $Cay(SL(2,\mathbb{Z}_{n});S_{n})$. This leads to the expression:
$H=\text{GNN}(\text{GNN}(X,A),A_{\text{Cay}(n)})$
The EGP model is purposefully designed to enhance connectivity for long-range tasks, thereby facilitating efficient communication between distant nodes.
Arnaiz et al. [59] introduced a novel rewiring methodology that synthesizes the concepts of commute time and graph spectral gap. They exploit the Lovász bound:
$\left|\frac{\text{Com}(u,v)}{\text{vol}(G)}-\left(\frac{1}{d_{u}}+\frac{1}{d_{v}}\right)\right|\leq\frac{1}{\lambda_{1}}\frac{2}{d_{\text{min}}},$ (9)
where $\lambda_{1}\geq 0$ represents the first non-zero eigenvalue of the normalized graph Laplacian matrix $\tilde{L}$. This methodology comprises two distinctive layers within a Graph Neural Network (GNN) based in each side of the inequality (9). The first layer, focuses on the left-hand side, the Commute Time Layer (CT-LAYER), serves as a differentiable, parameter-free component tailored to learn the commute time. Meanwhile, the second layer, focuses on the right-hand side, the Gap Layer (GAP-LAYER), functions as the rewiring layer, tasked with optimizing the spectral gap based on the specific network characteristics and task objectives. The CT-LAYER calculates the commute times between nodes and retains edges with significant effective resistance $R_{u,v}$ ($=\frac{\text{Com}(u,v)}{\text{vol}(G)}$) to preserve graph’s topology. Subsequently, the GAP-LAYER adjusts the graph’s adjacency matrix $A$ utilizing ratio-cut and normalized-cut approximations [68] to minimize the spectral gap. During training, both CT-LAYER and GAP-LAYER dynamically learn the weights to predict optimal topology modifications for unseen graphs during testing. This integrated framework endows the GNN with the capability to adaptively learn and implement rewiring strategies, effectively mitigating challenges associated with over-squashing while considering the intricacies of graph structure and task requirements. Banerjee et al. [46] introduced two graph rewiring techniques, namely Random Local Edge Flip (RLEF) and Greedy Random Local Edge Flip (G-RLEF), to alleviate bottlenecks in global information propagation within a graph. RLEF draws inspiration from the flip Markov chain [69, 70, 71], which transforms a connected graph into an expander graph with high probability. For G-RLEF, the authors utilized the relationship between effective resistance $R_{u,v}$ (Eq. 4) and the number of triangles, denoted by $\sharp_{\Delta}(u,v)=|N_{u}\cap N_{v}|$, that contain the edge $(u,v)$:
$R_{u,v}\leq\frac{2}{2+\sharp_{\Delta}(u,v)}$ (10)
They noted that a decrease in the number of triangles is accompanied by an increase in the spectral gap. Therefore, they devised a strategy to flip the edges $(u,i)$ and $(v,j)$ in a manner that minimizes the net change in the number of triangles. G-RLEF, a greedy variant of the RLEF algorithm, aims to expedite the spectral expansion process by employing a non-uniform sampling strategy for the hub edge $(u,v)$, where the selection is proportional to their effective resistance.
IV Unifying Approaches to Address Over-Squashing and Over-Smoothing Trade-offs in Graph Neural Networks
While these rewiring methods aim to enhance graph connectivity, they come with certain drawbacks, particularly when excessive modifications are made to the input graph. One prominent concern is the loss of valuable topological information inherent to the original graph. Additionally, the act of adding edges has a smoothing effect on the graph. If we introduce an excessive number of edges to the input graph, a standard Graph Convolutional Network (GCN) may encounter a common issue known as over-smoothing, as highlighted by Li et al. [72].
Over-smoothing in graph neural networks (GNNs) occurs when the embeddings of nodes from different classes become increasingly similar or indistinguishable [73, 72]. This phenomenon is particularly prevalent in multi-layer MPNNs designed for short-range tasks, where a node’s accurate prediction heavily relies on information from its immediate neighborhood. As the network’s depth increases, the embeddings tend to lose the ability to differentiate between nodes of different classes due to the repeated aggregation of local information. In contrast, over-squashing in graph neural networks (GNNs) occurs when the number of layers in the network increases to accommodate long-range tasks, the receptive field of each node expands to include more neighboring nodes [32, 35]. Consequently, the information gathered from these extended receptive fields is compressed into fixed-length node vectors, resulting in a loss of detailed information from distant nodes. This compression of information leads over-squashing, where the network fails to accurately convey messages originating from distant nodes.
Certain methodologies have emerged that tackle the intertwined challenges of over-smoothing and over-squashing in unison, establishing an interconnected relationship between these fundamental limitations within graph neural networks.
Nguyen et al in [40] reveals a connection between local graph geometry and the issues of over-smoothing and over-squashing and introduced a novel rewiring technique known as Batch Ollivier-Ricci Flow (BORF), which harnesses the power of Ollivier-Ricci curvature to address these interrelated challenges in GNNs. Mathematically over-smoothing can be seen as
$\sum_{(u,v)\in E}\left|h_{u}^{(k)}-h_{v}^{(k)}\right|\rightarrow 0\text{ as }k\rightarrow\infty$ (11)
Over-smoothing increases as number of layer $k$ increases as node features become indistinguishable. This definition is similar to the definition based on the node-wise Dirichlet energy
$\mathcal{E}(X)=\frac{1}{2}\sum_{i,j,k}A_{uv}\left(\frac{X_{uk}}{\sqrt{d_{u}}}-\frac{X_{vk}}{\sqrt{d_{v}}}\right)^{2}$ (12)
given in [74, 44] where $X$ is the feature matrix and $A$ is the adjacency matrix. In addition to the connection of edges with negative curvature as discussed in [35] also, Nguyen et al linked over-smoothing to edges with positive curvature. Specifically, they proved that, if every edge curvature in a regular graph $G$ is bounded from below by a sufficiently high constant then the difference between the features of any pair of neighboring nodes , exponentially converges to 00 in a typical GNN, i.e,
$\sum_{(u,v)\in E}\left|h_{u}^{(k)}-h_{v}^{(k)}\right|\leq c_{1}e^{-c_{2}k},$ (13)
where $c_{1},c_{2}$ are positive constant. BORF operates in batches and begins by identifying two sets of edges in each batch: $p$ edges with minimal curvature and $q$ edges with maximal curvature. It then optimizes the graph’s connectivity by adding connections to the minimally curved edges, ensuring efficient communication between distant nodes. This alleviates over-squashing. Additionally, BORF removes the maximally curved edges to prevent over-smoothing, as these can lead to excessive smoothing of node features. Furthermore, the algorithm’s flexibility allows it to operate as a net edge addition, subtraction, or net-zero rewiring, providing adaptability to different data characteristics. In the similar line, Liu et al. [45] connected over-smoothing with edges with positive curvature and over-squashing with negatively curved edges and addressed both problems by focusing on edge removal based on curvature metrics. They harnessed this insight to address the challenges of over-smoothing and over-squashing by introducing a sampling layer driven by Ricci curvature. This sampling layer selectively drops a portion of edges with low Ricci curvature at each GNN layer, effectively mitigating the issues associated with over-smoothing and over-squashing. Recently, a curvature-based edge dropping algorithm known as Curvature-Based Edge Dropping (CBED) is introduced in the work by Dai Shi et al. [75]. This innovative approach strategically removes edges with the highest positive curvature. By doing so, it aims to enhance the model’s adaptability to graphs characterized by heterophily and, in the process, alleviate the issue of over-smoothing.
Giraldo et al. [33] established a profound trade-off between over-smoothing and over-squashing. In the context where $P=D^{-1}A$ is the random walk transition matrix, for any initial distribution $f:V\rightarrow\mathbb{R}$ with $\sum_{v\in V}f(v)=1$, the distribution after $k$ steps is given by $f^{T}P^{k}$, where $f\in\mathbb{R}^{N\times 1}$ is the vector of initial distributions such that $f(i)$ represents the function evaluated on the $i$-th node. The random walk is ergodic when there is a unique stationary distribution $\pi$ satisfying $\lim_{s\rightarrow\infty}f^{T}P^{s}=\pi$. They proved the following results
Theorem IV.1
Let $h(G)$ be the Cheeger constant of $G$, and let $s$ be the number of required steps such that the $\ell_{2}$ distance between $f^{T}P^{s}$ and $\pi$ is at most $\varepsilon$. Then
$2h(G)\geq\frac{1}{s}\log\frac{{\max_{u\in V}\sqrt{d_{u}}}}{{\varepsilon\min_{v\in V}\sqrt{d_{v}}}}$
That is, if $s$ approaches 0, then $h_{G}$ approaches infinity, implying that we can reduce the bottleneck effect in the graph by speeding up the convergence to the stationary distribution. Conversely, if $h_{G}$ approaches 0, then $s$ approaches infinity, indicating that we can avoid converging to the stationary distribution by promoting a bottleneck-like structure in the graph. In response to this challenge, Giraldo et al. introduced the Stochastic Jost and Liu Curvature Rewiring (SJLR) algorithm, a notable departure from previous curvature-based techniques [35, 46, 44]. SJLR dynamically adds and removes edges during the training phase of GNNs while maintaining the fundamental graph structure unaltered during the testing phase. This adaptability sets SJLR apart as a promising approach to address the intricate challenges posed by over-smoothing and over-squashing in GNNs. To tackle the trade-off between these challenges, Karhadkar et al. [44] proposed a novel rewiring method called First-order Spectral Rewiring (FoSR) with the objective of optimizing the spectral gap of the graph input to the GNN.
Theorem IV.2
The first-order change in the second eigenvalue $\lambda_{1}=\lambda_{1}(D^{-1/2}AD^{-1/2})$ resulting from adding the edge $(u,v)$ is approximately equal to [44]:
$\frac{2\mu(u)\mu(v)}{(\sqrt{1+d_{u}})(\sqrt{1+d_{v}})},$ (14)
where $\mu$ represents the second eigenvector of $D^{-1/2}AD^{-1/2}$, and $\mu(u)$ denotes the $u$-th entry of $\mu$.
The FoSR algorithm carefully calculates the first-order change in the spectral gap caused by adding each edge and then chooses the edge that maximizes this change. In other words, it approximates $\mu$ and selects an edge that minimizes the expression given in Equation (14). Within this framework, the authors propose a comprehensive approach that not only introduces this innovative rewiring method but also integrates a relational Graph Neural Network to effectively leverage these rewired edges. They have demonstrated that the Dirichlet energy (12), consequently the over-smoothing in the proposed GNN is lower than that of traditional GNNs. Beaini et al. in [60] have associated these challenges with the incapacity of Graph Neural Networks (GNNs) to effectively capture directional information within graphs. This limitation constrains their ability to understand graph structures and perform feature transformations. To address this, they introduced Directional Graph Networks (DGNs), leveraging the eigenvectors $\lambda_{j}$ of the normalized Laplacian matrix $\tilde{L}$. They demonstrated that by propagating information in the direction of $\mu$, DGNs efficiently facilitate information sharing between distant nodes in the graph, thereby reducing the diffusion distance between them.
Theorem IV.3
Let $u$ and $v$ be two nodes in graph $G$ such that $\mu(u)<\mu(v)$. If $u_{0}$ is the node obtained by moving one step from $u$ in the direction of $\nabla\mu$, then there exists a constant $C$ such that for $C\leq t$, the diffusion distance
$d_{t}(u_{0},v)<d_{t}(u,v).$
This reduction in distance is proportional to $e^{-\lambda_{1}}$.
The diffusion distance at time $t$ between nodes $u$ and $v$ is given by:
$d_{t}(u,v)=\left(\sum_{u^{\prime}\in V}\left(p_{t}(u,u^{\prime})-p_{t}(u^{\prime},v)\right)^{2}\right)^{\frac{1}{2}},$
where $p_{t}(u,v)=P(P^{t}=v|u_{0}=u)$. In summary, the DGN model, through its globally consistent directional information, effectively addresses challenges such as over-squashing and over-smoothing. It empowers GNNs to comprehend local graph structures, perform meaningful feature transformations, and mitigate the adverse effects of these issues.
V Graph Transformers and other Strategies
Graph transformers have gained substantial attention as an alternative approach to combating over-smoothing and over-squashing in the context of graph and computer vision domains [76, 77, 78]. This approach leverages the inherent strengths of transformer architectures. for instance, Ying et al. [79] observed that transformers are less susceptible to over-smoothing compared to traditional GNNs. Their ability to model graph data efficiently contributes to mitigating the over-smoothing problem. Kreuzer et al. [80] highlighted the resilience of transformers to over-squashing. Transformers establish direct paths connecting distant nodes, which alleviates the over-squashing challenge.
However, it’s worth noting that transformers have limitations, including significant computational and memory requirements due to the need for every node to attend to all others. This can make them less suitable for large-scale graph applications and may result in improper training leading to a blend of local and non-local interactions. To tackle this challenge several methods are proposed.
TABLE II: Dataset Statistics
Node classification
Cora Citeseer Pubmed Film TwitchDE Tolokers Cornell Texas Wisconsin Chemeleon Squirrel Actor
#nodes 2708 3327 19717 7600 9496 11758 183 183 251 2277 5201 7600
#edges 10556 9104 88648 53504 153138 519000 280 295 466 31421 198493 26752
#features 1433 3703 500 932 2514 10 1703 1703 1703 2089 2089 932
#classes 7 6 3 5 2 2 5 5 5 5 5 5
$H(G)$ 0.81 0.74 0.80 0.22 0.63 0.59 0.30 0.11 0.21 0.23 0.22 0.22
Directed $\times$ $\times$ $\times$
Graph classification
NCI-1 NCI-109 Reddit-B Reddit-5K Reddit-12K Collab Enzymes BZR MUTAG PTC COX2 Proteins
#graphs 4110 4127 200 4999 11929 5000 600 405 188 344 467 1113
Avg. nodes 30 30 430 509 391 75 32.63 35.75 17.93 25.56 41.22 39.06
Avg. edges 32 32 498 595 457 2458 62.14 38.36 19.79 25.96 43.44 72.82
#features 37 38 0 0 0 0 0 0 0 0 0 1
#classes 2 2 2 5 11 3 6 2 2 2 2 2
Directed $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[…]
Definition V.1
For a given graph $G$, the Reaching Coefficient (RC) measures the average length of the shortest path from unlabeled nodes to their corresponding labeled nodes within their respective classes:
$RC=\frac{1}{|V^{U}|}\sum_{u\in V^{U}}\frac{1}{|V^{L}|}\sum_{v\in V^{L}}\left(1-\frac{\log d_{G}}{Diam(G)}\right),$
where $V_{L}$ and $V^{U}$ denotes labeled and unlabeled nodes.
The reaching coefficient indicates the extent of information propagation from labeled to unlabeled nodes by GNNs. It’s important to note that $RC\in[0,1)$, where a higher $RC$ signifies improved reachability
Definition V.2
For a graph $G$, the Squashing Coefficient (SC) quantifies the average Ricci curvature of edges along the shortest path from unlabeled nodes to the labeled nodes within their corresponding classes:
$SC=\frac{1}{|V^{U}|}\sum_{u\in V^{U}}\frac{1}{|N_{u}^{L}|}\sum_{v\in N_{u}^{L}}\frac{\sum_{(u^{\prime},v^{\prime})\in P_{uv}}Ric(u^{\prime},v^{\prime})}{d_{G}},$
where $N_{u}^{L}$ is the labeled neighborhood of node $u$, and $Ric(u,v)$ denotes the Ollivier-Ricci curvature [82].
The squashing coefficient may exhibit either positive or negative values, with a larger $SC$ indicating reduced squashing. PASTEL is designed to improve intra-class connectivity among nodes in GNNs by optimizing information propagation paths. This is achieved through the utilization of transformer based position encoding mechanism that captures the relative positions of unlabeled nodes with respect to labeled nodes. Furthermore, a class-wise conflict measure, employing Group PageRank, assesses the influence of labeled nodes from different classes, guiding adjustments to edge weights to enhance intra-class connectivity. Rongqin et al. [64] delve into the challenge of over-squashing within GNNs, linking it to message redundancy during aggregation. They establish that redundancy significantly contributes to over-squashing, especially when conventional GNNs struggle with long-length path propagation, limiting their ability to handle long-range interactions. To address this, they introduce the Redundancy-Free Graph Neural Network (RFGNN), leveraging extended paths (epaths) [^3] A path that has no repeated node except for the first node as the end node in a path of length more than 2., to capture complex graph structures. They implement truncated ePaths trees (TPTs) for message-passing, where TPTs of height $k$ ($TPT_{(G,u)^{k}}$) are obtained by running a BFS from node $u$ in graph $G$, accessing epaths of length up to $k$. RFGNN employs a path-search-tree concept constructed via breadth-first search to eliminate redundancy in message propagation, ensuring efficient information transmission without over-squashing. This novel de-redundancy technique balances epath influence, enhancing GNNs’ ability to capture structural information in original graphs while mitigating over-squashing. Tortorella and Mechelli [58] address the issue of over-squashing in node classification tasks within graphs characterized by low homophily (as defined in Equation (20)). They propose a reservoir computing model called the Graph Echo State Network (GESN).
Definition V.3
Reservoir computing [83] presents a paradigm for designing recurrent neural networks (RNNs) with efficiency. Input data is encoded by a randomly initialized reservoir, with only the readout layer requiring training for downstream task predictions.
GESN extends the reservoir computing paradigm to graph-structured data. Node embeddings are computed recursively through a nonlinear dynamical system represented by the equation:
$h^{(k)}_{u}=\text{tanh}\left(W_{\text{in}}x_{u}+\sum_{v\in N(u)}W_{\hat{h}}h^{(k-1)}_{v}\right),\quad h^{(0)}_{v}=0.$ (15)
Here, $W_{\text{in}}\in\mathbb{R}^{n\times d_{0}}$ and $W_{\hat{h}}\in\mathbb{R}^{n\times n}$ are the input-to-reservoir and recurrent weights, respectively, for a reservoir with $n$ units. Equation (15) iterates over $k$ until the system state converges to the fixed point $h^{(\infty)}_{u}$, which serves as the embedding. For node classification tasks, a linear readout is applied to node embeddings. The training-free characteristic makes GESN an efficient and effective solution for node classification tasks, offering a promising approach to mitigate issues of long-range message passing and over-squashing in heterophilic graphs.
Gravina et al. in [62] introduced the Anti-Symmetric Deep Graph Network (A-DGN), an innovative framework tailored to address the challenge of long-term information propagation in Deep Graph Networks (DGNs). This approach leverages principles from ordinary differential equations (ODEs) by representing node features as a function of time, $x_{u}(t)$ for $u\in V$, and defining a node-wise ODE as:
$\frac{dx_{u}(t)}{dt}=f_{G}(x_{u}(t)),$ (16)
for $t\in[0,T]$, with initial conditions $x_{u}(0)=x_{u}\in\mathbb{R}^{d_{0}}$. Here, $f_{G}:\mathbb{R}^{d_{0}}\rightarrow\mathbb{R}^{d_{0}}$ represents the dynamics of node representations. This ODE process can be seen as continuous information processing over graphs, starting from initial features $x_{u}(0)$ and culminating in final node features $x_{u}(T)$. This process shares similarities with standard DGNs, as it computes nodes’ states that can be used as an embedded representation of the graph and then fed into a readout layer for downstream tasks on graphs. Gravina et al. establish theoretical conditions under which a stable and non-dissipative ODE system can be realized on graph structures, utilizing anti-symmetric [^4] A square matrix $A$ is anti-symmetric if $A^{T}=-A$. weight matrices. The A-DGN layer is formulated through the forward Euler discretization of the obtained graph ODE:
$x_{u}^{\ell}=x_{u}^{\ell-1}+\epsilon UP\left((W-W^{T}-\gamma I)x_{u}^{\ell-1}\right.\\ \left.+AGG(x_{v}(t),\text{ where }v\in N_{u})\right),$ (17)
where $\gamma$ controls the strength of diffusion and $\epsilon$ is the discretization step. This process enforces specific properties on the ODE system, preserving long-term dependencies between nodes within the graph and alleviating the problem of over-squashing in GNNs. In their paper [61], Shao et al. introduced a novel algorithm, the Multi-Scaled Heat Kernel based Graph Neural Network (MHKG), aiming to address the challenge of over-smoothing as well as over-squashing in GNNs. They propose a generalized graph heat equation represented as:
$\frac{dH^{(t)}}{dt}=-f(\tilde{L})H^{(t)},$ (18)
with the initial condition $H^{(0)}=X\in\mathbb{R}^{n\times d_{0}}$, where $f(\tilde{L})$ acts as a filtering function operating element-wise on the eigenvalues of $\tilde{L}$, typically represented by polynomial or analytic functions. Mathematically, this is denoted as $f(\tilde{L})=Uf(\Lambda)U^{T}$, where $U$ is the matrix of eigenvectors and $\Lambda$ is the diagonal matrix of eigenvalues of $\tilde{L}$. Similar to the basic heat equation, they define $K_{t}=e^{-tf(\tilde{L})}$ as the generalized heat kernel. The ODEs (18) used in GNNs mimic heat flow in a fixed direction and speed, akin to the second law of thermodynamics, leading to over-smoothing, where all node features become equal over time. The proposed MHKG model introduces a mixed dynamic approach, combining both smoothing and sharpening effects on node features, formulated as:
$H^{(\ell)}=U\text{diag}(\theta_{1})\Lambda_{1}U^{\top}H^{(\ell-1)}W^{(\ell-1)}\\ +U\text{diag}(\theta_{2})\Lambda_{2}U^{\top}H^{(\ell-1)}W^{(\ell-1)},$ (19)
where $\text{diag}(\theta_{1})$ and $\text{diag}(\theta_{2})$ are trainable filtering matrices, and $\Lambda_{1}=-f(\Lambda)=\text{diag}{e^{-f(\lambda_{i})}}{i=1}^{N}$ and $\Lambda_{2}=f(\Lambda)=\text{diag}{e^{f(\lambda_{i})}}{i=1}^{N}$. The model manipulates the graph spectral domain through controlled adjustments of time, effectively enhancing node feature sharpness while managing the trade-off. Curvature based pooling technique is also proposed to deal with over-smoothing and over-squashing in GNNs. Curvature-based Pooling within Graph Neural Networks (CurvePool) [43] relies on the Balanced Forman curvature [35] to identify critical structures in the graph that contribute to these problems. This method calculates curvature values for each edge and employs a criterion to group nodes into clusters, ensuring that nodes with similar curvature profiles are pooled together. The resulting node clusters are transformed into new nodes in the pooled graph, and node representations within each cluster are aggregated using operators like mean, sum, or maximum. To retain the original graph structure, CurvPool remaps old edges to the new node clusters. By leveraging graph curvature to guide the pooling process, CurvPool effectively balances over-smoothing and over-squashing, ultimately improving the performance of GNNs in graph classification tasks.
Each of these approaches offers a unique perspective and set of techniques to address the challenges of over-squashing in graph-based machine learning models.
VI Datasets
The common datasets employed for node and graph classification tasks in the models listed in Table I are presented in Table II, along with detailed dataset statistics. It’s important to note that this list is not exhaustive, as there are numerous other datasets, including synthetic and large-scale real-world ones, utilized for various research purposes. Table II displays the statistics of the datasets used in this study, where $H(G)$ represents the graph’s homophily, as defined in [84], calculated as
$H(G)=\frac{1}{|V|}\sum_{v\in V}\dfrac{\text{$v$'s neighbors with the same label as $v$}}{N_{v}}$ (20)
For node classification tasks, we employ a diverse set of 12 datasets, encompassing graphs of varying sizes and characteristics.
Cora [85], CiteSeer [86], and PubMed [87] are examples of paper citation networks. In these datasets, node features are represented as bag-of-words extracted from paper content, and the goal is to classify research topics. Notably, these datasets exhibit high homophily. In contrast, the Film dataset is created based on actor co-occurrences on Wikipedia pages and is categorized into five groups. This dataset poses a node classification task with low homophily characteristics. TwitchDE, on the other hand, is a social network comprising German gamer accounts from Twitch, categorized as suitable for work or adult profiles. The classification task involves profiling these accounts. The Tolokers dataset represents a collaboration network derived from the crowdsourcing platform Toloka. The objective here is to determine user activity, considering the challenge of class imbalance, with the evaluation metric being the area under the ROC curve. Cornell, Texas, Wisconsin are additional node classification tasks, each originating from university-related interactions. The Cornell dataset comprises research paper citation data. The Texas dataset represents friendships in a Texas college, and the Wisconsin dataset is derived from a university-related network. Node features and specific targets for these datasets can vary. Chameleon[88], Squirrel[88], Actor [89] are also novel datasets introduced for node classification. Chameleon captures interactions within a university community. Squirrel is a network of interactions among squirrels in a park. The Actor dataset models collaborations among actors in the film industry. Each of these datasets presents unique characteristics and classification tasks.
For graph classification tasks, we utilize the following datasets: NCI-1 and NCI-109 datasets involve classifying molecules as cancerous or non-cancerous. The node input features are represented as one-hot encodings of atom types, while edges signify chemical bonds. In datasets like Reddit-B, Reddit-5K, and Reddit-12K, interactions between users in Reddit discussion threads are captured. The primary task associated with these datasets is to determine the type of subreddit to which a discussion belongs. Collab comprises ego-networks from three distinct scientific collaboration fields. Unlike the previous datasets, Reddit tasks, and Collab, these datasets do not have node input features. Enzymes is a bioinformatics dataset for graph classification. It involves classifying enzymes based on their structures and functions. The BZR dataset is a small molecule dataset used for graph classification tasks. It is commonly employed for evaluating graph-based machine learning algorithms. MUTAG is another bioinformatics dataset for graph classification, primarily used for evaluating chemical informatics algorithms. The task is to predict mutagenicity. PTC is a bioinformatics dataset for graph classification, focusing on carcinogenicity prediction. The graphs represent chemical compounds. COX2 is a small molecule dataset, often used to assess graph-based machine learning models in chemistry-related tasks. The classification task is centered around predicting the inhibition of the COX-2 enzyme. Proteins is a bioinformatics dataset used for graph classification. The task is to classify proteins based on their functions. These datasets are from Tudataset [90].
In all these tasks, we intentionally avoid introducing structural input features such as node degrees or positional encodings. A summary of relevant dataset statistics is provided in Table II for reference.
VII Conclusion and Future Direction
This survey has delved into the depths of over-squashing, unearthing its origins in information compression across distant nodes. The journey traversed a diverse array of strategies aimed at mitigating its impact – from innovative graph rewiring methods and curvature-based approaches to spectral techniques and the promise of graph transformers. As we tread this path, a nuanced interplay between over-smoothing and over-squashing has come into focus, demanding a balanced resolution. This exploration stands as a testament to the ongoing dialogue among researchers, driven by the pursuit of more refined and capable Graph Neural Networks. In closing, the quest to unravel over-squashing continues to be a beacon guiding our pursuit of more effective models, propelled by the dynamic nature of graph data.
As discussed in this survey, while techniques exist to mitigate the effects of over-squashing in GNNs, several directions remain open for exploration:
  1. 1.
    Dynamic Graph Adaptation: Existing models and rewiring methods in the literature are primarily designed for static graphs, overlooking the dynamic nature of real-world graph data. Investigating the performance of these techniques, as well as developing novel approaches, in the realm of dynamic graphs is essential.
  2. 2.
    Handling Distributional Shifts: Graph representation learning often struggles with generalizing to out-of-distribution data, leading to performance degradation in real-world scenarios. While techniques exist to mitigate distributional shifts in graph data, the impact of over-squashing in such contexts remains underexplored.
  3. 3.
    Robustness to Noisy Data: Investigating techniques to improve the robustness of GNNs to noisy or adversarial data, which can exacerbate issues such as over-squashing. This may involve exploring methods for robust training or designing inherently resilient models to perturbations in the input data.
  4. 4.
    Scalability and Efficiency: Addressing scalability and efficiency concerns associated with existing over-squashing mitigation techniques, especially for large-scale graph datasets. This may entail developing scalable algorithms or designing lightweight model architectures suitable for deployment in resource-constrained environments.
  5. 5.
    Incorporating Heterophily in Graph Classification: While discussed models incorporate heterophily only in node classification tasks, exploring its impact on graph classification tasks would be intriguing. Analyzing the role of heterophily in graph classification could lead to insights into improving the performance and robustness of graph classification models.
Acknowledgment
I extend my heartfelt appreciation to Dr. Karmvir Singh Phogat for providing invaluable insights and essential feedback on the research problem explored in this article. His thoughtful comments significantly enriched the quality and lucidity of this study.